OpenStudy (anonymous):

Still looking for the inverse function of y=x/(x^2-x-2) ... Thanks =]

OpenStudy (anonymous):

Hmm, it doesn't have inverse actually unless you restrict the domain.

OpenStudy (jamesj):

(For the record, you really shouldn't spam this question to @Physics.) Yes, you do need to think carefully about the domains and ranges. But as a first step, you can find the naive inverse by solving your equation for x. You can do that, can't you? Just work through the algebra. I.e., if y=x/(x^2-x-2) then yx^2 - yx - 2y -x = 0 i.e., yx^2 -(y+1)x - 2y = 0 Now use the quadratic formula.

OpenStudy (anonymous):

@jamesj i think we got exactly that before. quadratic formula may have been the issue from there on

OpenStudy (jamesj):

(@sat73, ah, I hadn't seen the earlier version.)

OpenStudy (anonymous):

maybe that was someone else. exact same question though.

OpenStudy (anonymous):

thanks for the reply, but i sill can't get it though. I've noticed that the range cannot be for x=2 and x=-1, but how to use the quadratic formula from this step exactly ?

OpenStudy (jamesj):

If yx^2 -(y+1)x - 2y = 0 then in standard form --- ax^2 + bx + c = 0 --- a = y, b = -(y+1), c = -2y.

OpenStudy (anonymous):

put \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] with \[a=y,b=-(y+1), c=2y\]

OpenStudy (anonymous):

lol \[c=-2y\] what jamesj said

OpenStudy (anonymous):

yes i did that but then i got \[\Delta\] which i don't know how to make sqrt out of it

OpenStudy (jamesj):

That's fine, it is what it is.

OpenStudy (anonymous):

and the plus, minus business comes from that fact that your original function is not one to one